### v2

# Political Interventions to Reduce Single-Use Plastics (SUPs) and Price Effects: An Event Study for Austria and Germany

Abstract
Single-use plastics (SUPs) impose substantial environmental costs. Following Directive (EU) 2019/904, Austria and Germany introduced producer charges and fund payments to finance clean-up. Using a high-frequency panel of retail offer spells with prices and a fixed-effects event-study design with two-way clustered standard errors, this paper estimates the extent to which these costs are passed through to consumer prices. We find clear evidence of price pass-through in Austria. Pooled Austrian SUP products are 13.01 index points more expensive than non-SUP controls within twelve months (DiD(12m); p<0.001p{<}0.001) and 19.42 points over the full post-policy period (p<0.001p{<}0.001). At the product level, highly taxed balloons exhibit strong and persistent effects (DiD(12m)=13.43, p=0.007p{=}0.007; Full DiD=19.96, p<0.001p{<}0.001). For plastic to-go cups, the twelve-month estimate is negative but statistically insignificant (DiD(12m)=-22.73, p=0.096p{=}0.096), while the full-period estimate is positive and likewise insignificant. In Germany, where the Single-Use Plastics Fund took effect in 2024, the post-policy window is short and estimates are not statistically significant; these results are therefore interpreted as descriptive rather than causal. As the data contain prices but not quantities, the analysis speaks to price incidence on consumers and producers, not to changes in consumption or litter.
Keywords: single-use plastics; environmental policy; extended producer responsibility; deposit–refund; price pass-through; event study; difference-in-differences; retail prices; Austria; Germany.
JEL codes: Q53; Q58; H23; C23; D12.
1 Introduction
On EU beaches, plastics account for 80–85% of litter items; about half of those plastic items are single-use plastics (EU,, 2019). Cities and towns carry much of this load. Street sweeping and waste collection tied to plastic trash run into billions of euros each year, costs often shouldered by the public purse (EU,, 2019). Environmental policy redirects part of these costs back to producers and consumers through taxes, fees, and producer responsibility so that market prices come closer to social costs (Fullerton and Metcalf,, 2001). Under imperfect competition, the degree of price pass-through depends on demand curvature, cost shares, and market conduct (Weyl and Fabinger,, 2013). Evidence from excise and sales taxes shows that pass-through to retail prices is often substantial (Marion and Muehlegger,, 2011; Besley and Rosen,, 1999), and small levies on plastic bags have produced large consumption responses (Convery et al.,, 2007; Homonoff,, 2018). Deposit–refund systems form a related design used to curb litter and improve collection (Walls,, 2011).
The sample period from 2020 to 2025 features several overlapping shocks. The COVID-19 pandemic and its recovery phase generated broad supply and demand shifts and high inflation. From early 2022, the war in Ukraine added a further energy and commodity price shock that raised costs for plastics and packaging producers (Packaging Europe,, 2022; Plastics Europe,, 2022). This macro context matters for identification, since any attempt to isolate the effect of SUP charges must be identified inside this turbulent time window.
Member states in the EU moved on several fronts after Directive (EU) 2019/904: bans for selected items, labels, and producer responsibility rules (EU,, 2019). Austria and Germany moved to producer charges and fund payments that finance clean-up and reporting. This paper asks a narrow question inside that broader effort: do these upstream charges reach the shelf? If so, relative prices should move in favor of non-SUP alternatives.
Methodologically, the paper falls within the class of event-study and difference-in-differences designs used to analyze staggered policy adoption and dynamic, heterogeneous treatment effects (Bertrand et al.,, 2004; Callaway and Sant’Anna,, 2021; Sun and Abraham,, 2021). The empirical setting is a retailer–product panel with relatively short price spells and a very broad product tree spanning categories, subcategories, and sub-subcategories. The empirical strategy relies on a parsimonious fixed-effects event-study framework and shifts the main causal interpretation to (i) within-country comparisons between SUP and non-SUP products and (ii) a careful discussion of remaining confounders and omitted variables.
We use monthly price data for items subject to SUP regulation and for comparator products outside the regulation in Austria and Germany, drawing comparators from the full set of available offer spells rather than restricting them to the same product space, and assuming neither substitution nor complementarity. The empirical strategy is an event-study design with product and retailer fixed effects and two-way clustered standard errors.
2 Policy Background
Background is the EU Directive (EU) 2019/904, which aims to cut the damage from specific plastic products by pushing producers and buyers to face clean-up costs and by reducing items that are hard to collect (EU,, 2019). Member states received latitude on the exact toolkit.
2.1 Austria: Producer Charges and Reporting (Treatment Year 2022)
Austria relies on producer charges tied to quantities sold and a digital reporting portal. Since 2023, producers report volumes through ARA Online. In 2022, Austria added product-based charges reflecting cleanup costs:
  • Food packaging, beverage cups, plastic wraps: €225 per ton
  • Tobacco products with plastic filters: €450 per ton
Micro-enterprises or very small quantities can opt for a €13 flat payment (BMK,, 2024). Revenues support local waste crews, litter prevention, and upgrades to collection and sorting.
2.2 Germany: Single-Use Plastics Fund Act (EWKFondsG, 2024)
Germany runs a central fund. The Einwegkunststofffondsgesetz took effect on January 1, 2024. Producers register in DIVID and pay per product and weight. The Federal Environment Agency (UBA) manages inflows and distributes money to municipalities for street cleaning, bin maintenance, and disposal linked to SUPs (UBA,, 2024). Item-specific fees mirror documented cleanup burdens.
Because the German charge starts only in 2024 and our sample ends in mid-2025, the post-policy period for Germany is short. This limits what can be learned about longer-run price adjustment in that country.
3 Data and Methodology
Data coverage.
The working dataset is a monthly product–retailer panel spanning January 2020 to May 2025 built from high-frequency offer spells. For each product $i$ and retailer $j$, we observe prices in month $t$. Items include products treated under SUP rules (e.g., balloons, to-go plastic cups) and non-SUP controls. Precise construction and control selection appear in Appendix A.
Product space and data spells.
The underlying database covers a very large catalog. Products sit in a deep tree of categories, subcategories, and sub-subcategories. Individual items enter and leave the catalog at different points, so price spells vary in length and timing. A keyword-based product search across this tree builds the SUP set and the control pool in a way that already compresses this complexity. Any alternative sampling of products for extended robustness checks or additional control groups would require a new pass through this tree and new filtering of data spells. That step is technically feasible, but costly and sensitive to classification choices, and lies beyond the present empirical work.
Aggregation and index.
Raw quotes are aggregated to product–retailer–month $(i,j,t)$. Let $\overline{p}_{ijt}$ be the average price in month $t$ and define the base month as February 2022 ($t=0$). We form a simple interpretable product–retailer specific index
$P_{ijt}\equiv 100\times\frac{\overline{p}_{ijt}}{\overline{p}_{ij0}}\,,\qquad t\in[-24,36].$
Our preferred outcome is the level index $Y_{ijt}=P_{ijt}$ (percent of the February 2022 level). Reading the index is straightforward: a value of $115$ means that prices are 15% higher than in February 2022. Event time is $e_{t}=t-0=t$ and we bin to 3-month steps $b\in\{\ldots,-6,-3,0,3,6,\ldots\}$ via floored $\text{bin}_{t}=\lfloor e_{t}/3\rfloor\cdot 3$ with bin 0 as the reference.
Alternative specifications in logarithms or first differences would imply a different scaling of coefficients and could affect small-sample precision. More importantly, because the data consist of retail offer spells of heterogeneous and often short duration, such transformations would mechanically exclude single-period spells and amplify noise in short sequences, thereby changing the effective estimation sample. Estimating these variants would therefore require additional sampling and cleaning of price spells to restore comparability, which is beyond the scope of this paper. The analysis consequently remains in levels, using the price index and interpreting effect sizes directly in index points.
Event-study specification.
For each group $g\in\{\text{SUP},\text{Non-SUP}\}$, we estimate the following event-study regression on the corresponding subsample; treated and control products are therefore estimated separately:
$Y_{ijt}=\sum_{b\in\mathcal{B}\setminus\{0\}}\beta^{(g)}_{b}\,\mathbf{1}\{\text{bin}_{t}=b\}\;+\;\alpha_{i}\;+\;\delta_{j}\;+\;\varepsilon_{ijt},$ (1)
where $\alpha_{i}$ denote product fixed effects, $\delta_{j}$ retailer fixed effects, and $\mathcal{B}$ is the set of three-month event-time bins spanning 24-24 to +36+36 months. The omitted reference category is bin 0 (February 2022), such that $\beta^{(g)}_{b}$ capture deviations from that baseline. Time fixed effects are deliberately excluded to preserve identifying variation from common shocks that may differentially affect SUP and non-SUP products within retailers. Standard errors are two-way clustered by product and retailer (Cameron et al.,, 2011).
The specification follows standard practice in fixed-effects event-study and difference-in-differences designs with dynamic treatment effects (Bertrand et al.,, 2004; Callaway and Sant’Anna,, 2021; Sun and Abraham,, 2021). Given the short and heterogeneous offer spells in the data and the absence of a clearly defined staggered adoption at the product level, the analysis remains within a parsimonious fixed-effects framework and emphasizes transparent within-country comparisons. Future research could build on this design by applying recently developed estimators for heterogeneous and staggered treatment effects—for example, grouping or interaction-weighted estimators (Callaway and Sant’Anna,, 2021; Sun and Abraham,, 2021)—once longer post-treatment panels or richer quantity data become available.
Common shocks and time indicators.
The period 2020–2025 is dominated by broad price swings linked to COVID-19, the war in Ukraine, and associated energy and commodity shocks (Packaging Europe,, 2022; Plastics Europe,, 2022). A natural reaction is to introduce calendar-time fixed effects (for example, month dummies, or country-by-month dummies) in addition to the event-study bins. In the present setup this choice comes with two issues.
First, the outcome is a product–retailer index normalized by February 2022. Much of the pure macro trend already appears in both SUP and non-SUP series and cancels in their difference. Time indicators would mainly soak up average price growth common to the full sample, at the cost of additional sampling variation in a panel where many product spells are short. Second, event-time dummies and rich time effects interact with the complicated product tree and the limited number of SUP items in some segments. A dense set of time interactions would require a fresh pass through the database and a more balanced sampling of products and retailers, which is difficult in the current data environment with scattered data spells.
For these reasons the main specification keeps product and retailer fixed effects, uses the index normalization at February 2022, and leans on a control design that pairs SUP and non-SUP items with similar retail channels. Common shocks such as COVID-19 and higher energy prices then appear in both treated and control series; the DiD contrasts measure relative movements. This strategy reduces the influence of macro shocks but does not erase them, which is why remaining confounding is discussed openly in the results and discussion sections.
Windowed DiD summaries.
For communication, we report “DiD” summaries as simple averages of post- and pre-event bin coefficients within symmetric windows and then take the difference:
$\mathrm{DiD}^{(g)}(w)\equiv\frac{1}{|\mathcal{B}^{+}_{w}|}\sum_{b\in\mathcal{B}^{+}_{w}}\beta^{(g)}_{b}\;-\;\frac{1}{|\mathcal{B}^{-}_{w}|}\sum_{b\in\mathcal{B}^{-}_{w}}\beta^{(g)}_{b}$,
$\mathcal{B}^{-}_{w}=\{-w,-w+3,\ldots,-3\},\qquad\mathcal{B}^{+}_{w}=\{3,6,\ldots,w\}.$
We report $w\in\{6,12\}$ months and a “Full” contrast averaging over all $b<0$ and all $b>0$. These are linear combinations of (1) coefficients; figures display 90% CIs for the individual bins, and we annotate DiD values for key subsamples.
Design choices, controls, and pre-trends.
Treated vs. control contrasts rely on proximity of products in function and use while guarding against substitution and sector-specific shocks. The main control pool consists of non-SUP goods that share retail channels and broad product context. In some specifications, a strict control set based on technologically distant products such as graphics cards is used as a stress test. These items share the same platform and digital environment but face markedly different supply and demand shocks; their purpose is to probe how far results survive under very different sectoral conditions, not to serve as the sole counterfactual path.
Pre-event bins offer a visual parallel-paths check: if treated and control series move in similar ways before the charge, the DiD interpretation looks more credible. The tables in Section 4 report selected event bins; patterns across bins show that pre-policy coefficients often deviate from zero in both treated and control groups but are relatively stable over time for given product–retailer pairs once normalized to February 2022. Some categories display modest pre-trends, so the strength of the identifying assumption varies by product.
Extended placebo exercises (for example, estimation on non-SUP products only with a fictitious treatment date, or alternative sets of control groups such as SUP-like products in non-treated countries) would deliver further reassurance. Constructing these variants with the present database would require fresh sampling across the deep product tree and a new round of keyword search, with careful handling of short data spells and category coverage. That step is left to future work, and the discussion section treats remaining confounding as a key limitation.
4 Results
This section documents dynamic price responses for Austrian balloons and to-go cups, pooled Austrian SUP products, and the combined Austria–Germany sample. Figures display the full sequence of three-month event-time bins. Tables list only those event bins for which at least one coefficient is statistically different from zero at the 10% level.
4.1 Summary of main DiD estimates
Table 1 summarizes the main difference-in-differences contrasts for Austria that meet conventional significance thresholds. All estimates are reported in index points relative to February 2022, so a value of +10+10 corresponds to a price level that is 10% higher than in the base month.
Table 1: Significant DiD contrasts for Austria (index points relative to February 2022)
Product / country Window DiD estimate p-value
Pooled SUP (Austria) 12 months post vs. pre 13.01 <0.001<0.001
Pooled SUP (Austria) Full post vs. pre 19.42 <0.001<0.001
Balloons (Austria) 12 months post vs. pre 13.43 0.007
Balloons (Austria) Full post vs. pre 19.96 <0.001<0.001
Cups (Austria) 12 months post vs. pre -22.73 0.096
  • Notes: Each row reports a windowed DiD contrast as defined in Section 3. Only coefficients significant at the 10% level are listed. Estimates are based on level indices with February 2022 as base month ($P=100$).
The pooled Austrian SUP sample and the balloon category display positive price responses relative to non-SUP controls. For to-go cups, the twelve-month contrast is negative and only marginally significant, pointing to greater volatility and weaker separation between treated and control products during the first year after the charge.
4.2 Balloons (Austria)
Figure 1: Balloons: SUP vs. Non-SUP (event study, 3-month bins). Notes: event time relative to February 2022; fixed effects by product and retailer; two-way clustered standard errors. Controls avoid direct substitutes to limit substitution and generalized inflation effects.
Before the introduction of the charge, both treated balloons and control products are priced below the February 2022 base level. Early pre-policy coefficients for treated items range between roughly 20-20 and 30-30 index points, while coefficients for controls are smaller in magnitude. Movements across pre-policy bins are gradual, with no visible shift immediately preceding the policy.
After the charge, prices for control products increase steadily and reach values between 10 and 25 index points in later bins. Treated balloon prices move back toward the base level and remain close to zero or slightly above. This divergence implies an increase in the relative price of balloons compared with controls. The twelve-month and full-period DiD contrasts in Table 1 summarize this pattern.
Table 2: Balloons: Treated vs. Control (Event-Study Coefficients, significant bins only)
Treated Control
Event bin Est. SE Est. SE
-24 -33.48 7.85
-21 -27.29 7.45
-18 -27.22 7.93
-15 -23.89 7.36
-12 -28.42 7.53 -3.92 1.76
-9 -26.30 7.26 -3.49 1.80
-3 -2.49 0.85
3 4.14 1.54
6 7.13 2.15
9 8.97 2.44
12 10.22 2.67
15 10.95 2.73
18 17.84 6.98
21 13.95 3.27
24 24.60 11.86
27 15.68 2.97
30 5.91 2.79 15.61 2.98
33 16.51 3.43
Observations 1,443 69,473
RMSE 19.00 146.20
Adj. R2 0.56 0.88
  • Notes: OLS with product and retailer fixed effects. SEs clustered by product and retailer. Only event bins with at least one coefficient significant at the 10% level are listed. Reference bin is 0 (February 2022).
The joint evidence from Figure 1 and Table 2 points to positive pass-through for Austrian balloons relative to controls, even during a period marked by broad price pressures affecting many sectors.
4.3 To-Go Cups (Austria)
Figure 2: Cups: SUP vs. Non-SUP (event study, 3-month bins). Notes: event time relative to February 2022; fixed effects by product and retailer; two-way clustered standard errors.
For to-go cups, both treated and control prices lie below the February 2022 base in the pre-policy period, with moderate variation across bins. Following the charge, both series display sizable short-run movements. The twelve-month DiD contrast in Table 1 is negative and marginally significant, reflecting the absence of a clear and sustained divergence between treated cups and controls during the first post-policy year once overall volatility is taken into account.
Over the full sample period, the DiD estimate becomes positive but remains imprecisely estimated. Evidence for pass-through in this segment is therefore weaker than for balloons or for the pooled Austrian SUP sample. Differences in product design, contracting arrangements, and competition from reusable alternatives may contribute to this pattern. Evaluating these channels would require additional data on costs and quantities, which are not available in the present dataset.
4.4 Austria: Within-Country SUP vs. Non-SUP
Figure 3: Austria: SUP vs. Non-SUP (event study, 3-month bins). Notes: event time relative to February 2022; fixed effects by product and retailer; two-way clustered standard errors. Controls avoid direct substitutes to reduce substitution and common-shock bias.
Pooling all Austrian SUP products provides a broader perspective across categories. Prior to the charge, both treated and control indices lie below the base month and increase gradually across early bins. Treated prices tend to sit somewhat below control prices during the pre-policy period, indicating level differences rather than a discrete change at the time of the policy.
After the charge, treated prices rise toward and above the base level, while control prices also increase but at a slower pace. In later bins, treated prices exceed February 2022 levels by several index points, with controls showing smaller deviations relative to their own pre-period averages. The windowed DiD contrasts in Table 1 capture these dynamics and point to economically meaningful pass-through for Austrian SUP products on average.
Table 3: Austria: Treated vs. Control (Event-Study Coefficients, significant bins only)
Treated Control
Event bin Est. SE Est. SE
-24 -25.75 7.12
-21 -21.80 6.06
-18 -22.71 6.52 -11.61 0.12
-15 -19.33 6.48 -12.03 0.11
-12 -22.26 6.84
-9 -21.53 6.13
-6 -9.39 0.36
3 -5.04 2.35
15 9.41 1.39
18 9.88 0.99
24 10.54 1.48
30 4.57 0.88
33 9.50 0.71
Observations 1,781 765
RMSE 18.40 5.02
Adj. R2 0.57 0.90
  • Notes: OLS with product and retailer fixed effects. SEs clustered by product and retailer. Only event bins with at least one coefficient significant at the 10% level are listed. Reference bin is 0 (February 2022).
4.5 Cross-Country Sample: SUP vs. Non-SUP
Figure 4: Cross-Country Sample (Austria and Germany): SUP vs. Non-SUP (event study, 3-month bins). Notes: event time relative to February 2022; fixed effects by product and retailer; two-way clustered standard errors. Controls avoid direct substitutes across countries as well.
The cross-country sample combines Austrian and German products and reflects differences in policy timing and design. Austria contributes observations from a charge in place from 2022 onward, while Germany enters with a fund that applies from 2024. Short-run fluctuations in Figure 4 are larger than in the Austria-only sample, and confidence intervals widen, reflecting both macroeconomic disturbances and the limited number of post-charge months for Germany.
When post-policy bins are averaged in the pooled cross-country sample, the contrast between SUP and non-SUP products is positive but imprecisely estimated. Country-specific splits, not reported in tabular form, indicate that this pattern is driven mainly by Austrian products. German SUP segments show smaller and statistically weak movements over the short window from 2024 to mid-2025. The German evidence should therefore be read as preliminary, with limited power to detect gradual price responses over longer horizons.
5 Discussion and Policy Implications
5.1 Price incidence and control groups
Clear price responses appear in Austrian retail data when producer charges are precisely defined and supported by established reporting systems (e.g. ARA Online in Austria and DIVID in Germany) (BMK,, 2024; UBA,, 2024). The estimated magnitudes for pooled SUP products and for balloons are consistent with prior evidence on taxes and charges applied to energy products and plastic bags (Marion and Muehlegger,, 2011; Besley and Rosen,, 1999; Convery et al.,, 2007; Homonoff,, 2018). In these earlier settings, relatively small per-unit charges translated into visible retail price adjustments and meaningful changes in economic incentives.
Control group construction plays a central role in this analysis. The primary control set consists of non-SUP products that share retail channels, digital platform exposure, and broad consumption environments with treated items. A separate strict control group based on products such as graphics cards contrasts SUP items with goods subject to very different upstream cost dynamics, including semiconductor supply cycles and cryptocurrency-driven demand. Estimates based on this strict set are best interpreted as a stress test for sector-specific pressures rather than as the main counterfactual path for SUP packaging. The substantive findings therefore rely on the broader non-SUP controls and on within-country comparisons.
The sample period coincides with substantial macroeconomic disturbances, including the COVID-19 pandemic and the war in Ukraine. Energy and feedstock prices rose sharply from 2022 onward, affecting plastics manufacturing costs across Europe (Packaging Europe,, 2022; Plastics Europe,, 2022). The empirical design addresses these forces through three features: (i) normalization of prices to a common base month; (ii) use of non-SUP controls that share retailers and platform-level exposure; and (iii) examination of pre-event bins for similar price movements. These steps do not rule out all sources of confounding, and part of the estimated price response may reflect differences in how SUP and non-SUP sectors react to shared cost shocks. This possibility informs the interpretation of the results.
5.2 Germany, the war in Ukraine, and macroeconomic context
For Germany, the post-policy window remains limited. The Single-Use Plastics Fund took effect in 2024, and the sample ends in May 2025, leaving only a small number of three-month event bins fully within the post-charge period. Over the same interval, German retail prices responded to war-related energy shocks and ongoing supply-chain adjustments affecting plastics and packaging costs (Packaging Europe,, 2022; Plastics Europe,, 2022). These concurrent forces complicate attempts to isolate the price effects of the fund in cross-country comparisons.
The German estimates are therefore presented as preliminary and descriptive. They provide an early view of price movements in treated segments relative to controls after the fund’s introduction but do not support precise quantitative statements about pass-through. A more complete assessment would require longer post-policy panels, broader coverage of German SUP products, and possibly additional macroeconomic series. Carrying out such an analysis would involve further sampling of the underlying database and the incorporation of external data sources, which falls outside the scope of the present study.
5.3 Quantities, welfare, and future research
The database used in this study contains detailed information on retail prices but does not observe quantities or sales volumes. The analysis therefore addresses only the first step in the policy transmission chain: whether producer charges and fund payments associated with single-use plastics (SUPs) are passed through to shelf prices. It does not address the subsequent steps—whether higher prices reduce SUP consumption, induce substitution toward reusable alternatives, or ultimately lower litter and environmental damages.
A simple CES model and testable predictions.
To clarify the key margins, consider a representative consumer who allocates expenditure between a SUP composite good $q_{S}$ and a reusable (or otherwise non-SUP) composite $q_{R}$ with CES preferences,
$U(q_{S},q_{R})=\left[\omega q_{S}^{\frac{\sigma-1}{\sigma}}+(1-\omega)q_{R}^{\frac{\sigma-1}{\sigma}}\right]^{\frac{\sigma}{\sigma-1}},\qquad\sigma>0,$ (2)
subject to the budget constraint $P_{S}q_{S}+P_{R}q_{R}=I$, where $P_{S}$ and $P_{R}$ are price indices and $I$ is income. The elasticity of substitution between SUP and non-SUP composites is $\sigma$. Standard CES demand implies
$q_{S}=\omega\left(\frac{P_{S}}{P}\right)^{-\sigma}\frac{I}{P_{S}},\qquad q_{R}=(1-\omega)\left(\frac{P_{R}}{P}\right)^{-\sigma}\frac{I}{P_{R}},$ (3)
where $P=\left[\omega P_{S}^{1-\sigma}+(1-\omega)P_{R}^{1-\sigma}\right]^{\frac{1}{1-\sigma}}$ is the CES price index.
On the supply side, let the pre-policy marginal cost of a representative SUP product be $c_{S}$, and suppose a per-unit producer charge $\tau$ increases marginal cost to $c_{S}+\tau$. Let $\rho\in[0,1]$ denote the pass-through rate into the consumer price (under competition, market power, and/or nominal rigidities), so that
$\Delta P_{S}=\rho\,\tau.$ (4)
The empirical analysis in this paper speaks directly to (4) by estimating changes in $P_{S}$ for SUP products relative to non-SUP products.
This simple structure yields three predictions that can be taken to data once quantities (or adequate proxies) are available:
  1. 1.
    Pass-through: a positive producer charge raises retail SUP prices, $\Delta P_{S}>0$, with magnitude governed by $\rho$.
  2. 2.
    Demand response and substitution: higher SUP prices reduce SUP demand according to the CES elasticity, with (semi-)elasticity
    $\frac{\partial\ln q_{S}}{\partial\ln P_{S}}=-\sigma-(1-\text{s}_{S})\sigma,\quad\text{and approximately}\quad\Delta\ln q_{S}\approx-\sigma\,\Delta\ln\!\left(\frac{P_{S}}{P_{R}}\right)\ \ \text{for small changes},$ (5)
    where $\text{s}_{S}$ is the expenditure share on $q_{S}$. [^1] The exact own-price elasticity under CES depends on how the aggregate price index $P$ moves with $P_{S}$. For empirically relevant small price changes, the approximation in relative prices is often a convenient sufficient statistic.
  3. 3.
    Environmental impact: if environmental damages are proportional to SUP quantity, $D=\delta q_{S}$ with $\delta>0$, then the policy reduces damages whenever it reduces $q_{S}$, with $\Delta D=\delta\,\Delta q_{S}<0$ if $\Delta q_{S}<0$.
Welfare accounting.
To connect the model to welfare, let total welfare be
$W=CS+\Pi+T-D,$ (6)
where $CS$ denotes consumer surplus, $\Pi$ producer profits, $T$ net public revenue from the charge (or fund payments), and $D$ environmental damages. A per-unit charge affects welfare through: (i) incidence (how much of $\tau$ is borne by consumers vs. producers, governed by $\rho$), (ii) the quantity response (governed by $\sigma$), and (iii) the conversion of reduced SUP use into environmental benefits (governed by $\delta$ and the mapping from SUP consumption to litter or waste).
In the CES setting, a convenient sufficient-statistics representation for the efficiency component of welfare change from a small tax increase is the Harberger triangle relative to the external benefit. For a marginal change $d\tau$,
$dW\approx\underbrace{\delta\,(-dq_{S})}_{\text{environmental benefit}}\;-\;\underbrace{\tfrac{1}{2}\,dq_{S}\,dP_{S}}_{\text{DWL from reduced consumption}},\qquad dP_{S}=\rho\,d\tau,$ (7)
where the demand slope (hence $dq_{S}$) is pinned down by $\sigma$ and baseline shares via (3)–(5). This expression highlights why price pass-through alone is not sufficient for welfare analysis: without $q_{S}$ (or credible proxies), one cannot quantify either the deadweight loss term or the environmental benefit term.
Implications for the present evidence.
The estimates in this paper provide evidence of positive pass-through ($\rho>0$) in Austria for several SUP segments, consistent with Prediction 1. However, Predictions 2–3 and the welfare decomposition in (7) require data on quantities and substitution patterns. In particular, even large price effects may translate into modest environmental benefits if $\sigma$ is small, whereas smaller price effects could yield substantial benefits if substitution away from SUPs is highly elastic.
Future research direction using modern DiD/event-study designs.
A natural next step is to construct linked panels that add (i) product-level sales volumes, or (ii) credible proxies for quantities (e.g. scanner data, retailer transaction aggregates, or platform-level order counts), and/or (iii) municipal outcomes such as litter counts, waste collection volumes, or recycling rates. With such data, the research designs emphasized in Bertrand et al., (2004); Callaway and Sant’Anna, (2021); Sun and Abraham, (2021) can be deployed to strengthen causal interpretation in three ways.
[…]
Implementing these extensions would require new sampling across the large product taxonomy, careful keyword design, and harmonized product classifications across countries. The present study prepares the ground by documenting a transparent pipeline and identification strategy and by clarifying where macroeconomic shocks and the absence of quantities currently limit welfare and environmental conclusions.
6 Conclusion
This paper studies price responses to producer charges targeting single-use plastics using high-frequency retail data from Austria and Germany. For Austria, the results show price pass-through both at the individual product level and in pooled analyses. Prices for balloons and for the pooled SUP group rise relative to non-SUP controls when measured as index values with February 2022 as the base month. For plastic to-go cups, price movements in the first year after the charge are more volatile, and while the full-period contrast is positive, it is estimated with limited precision.
For Germany, the evidence should be read as early and descriptive. The Single-Use Plastics Fund took effect only in 2024, and the sample ends in mid-2025, leaving a short post-policy window. This restricts the ability to assess longer-run price responses. In the pooled cross-country analysis, estimated effects mainly reflect the Austrian experience combined with a brief and variable German post-treatment period.
The empirical design relies on a fixed-effects event-study specification with three-month event-time bins and two-way clustered standard errors. The approach remains intentionally parsimonious and stays within the standard difference-in-differences framework, reflecting the structure of the available product–retailer panel and the uneven length of price spells. Extensions that allow for richer forms of treatment effect heterogeneity, alternative weighting schemes, or staggered adoption at finer levels would require longer post-policy panels and additional data. Once such data become available, recent advances in the difference-in-differences literature provide natural tools for further analysis (Bertrand et al.,, 2004; Callaway and Sant’Anna,, 2021; Sun and Abraham,, 2021).
Statements and Declarations
Funding: None.
Conflicts of interest: The author declares no conflicts of interest.
Data and code availability: Aggregated tables and replication code will be made available upon request; raw price microdata are subject to platform licensing.
Ethical approval: Not applicable.
Appendix A Data pipeline, sample construction, and control selection
This appendix documents the construction of the analysis dataset, the definition of treatment and control groups, and the aggregation and indexing procedures used in estimation.
A.1 Notation and relational operators
[…]
Helper functions.
$\texttt{unix\_to\_week}(ts),\quad\texttt{unix\_to\_month}(ts),\quad\texttt{months\_between}(x,y),$
$\texttt{sup}(name):\text{ TRUE if }name\ \texttt{ILIKE}\ p\text{ for some pattern }p\in\mathcal{P}.$
Base relations.
[…]
A.2 Cohort definition and event window
Let $T^{\ast}=1{,}577{,}836{,}800$ denote January 1, 2020 in Unix time. We restrict attention to products entering the database after this date:
$\textbf{Prod}^{+}\;:=\;\sigma_{\text{born\_ts}>T^{\ast}}(\textbf{Prod}).$ (8)
Time variables are attached to offers and clicks as follows:
$\textbf{Off}^{t}$ $=\mu_{\ \text{week}=\texttt{unix\_to\_week}(ts),\ \text{month}=\texttt{unix\_to\_month}(ts),\ \text{e}=\texttt{months\_between}(\text{month},m_{0})}(\textbf{Off}),$ (9)
$\textbf{Clk}^{t}$ $=\mu_{\ \text{week}=\texttt{unix\_to\_week}(ts),\ \text{month}=\texttt{unix\_to\_month}(ts),\ \text{e}=\texttt{months\_between}(\text{month},m_{0})}(\textbf{Clk}).$
We retain observations within the event window [24,36][-24,36] months and restrict to cohort products:
$\textbf{Off}^{win}$ $=\sigma_{-24\leq e\leq 36}(\textbf{Off}^{t})\;\bowtie_{\text{prod\_id}}\;\pi_{\text{prod\_id}}(\textbf{Prod}^{+}),$ (10)
$\textbf{Clk}^{win}$ $=\sigma_{-24\leq e\leq 36}(\textbf{Clk}^{t})\;\bowtie_{\text{prod\_id}}\;\pi_{\text{prod\_id}}(\textbf{Prod}^{+}).$
A.3 Retailer snapshot
Retailer attributes are fixed at their most recent observed values. Let $t^{\star}(\text{ret\_id})$ denote the last timestamp for a retailer:
$\textbf{Ret}^{max}$ $=\gamma_{\text{ret\_id}}\big(\max(ts)\rightarrow t^{\star}\big)(\textbf{Ret}),$ (11)
$\textbf{Ret}^{snap}$ $=\pi_{\text{ret\_id},\ ret\_name,\ \ldots}\Big(\textbf{Ret}^{max}\;\bowtie_{\textbf{Ret}.ret\_id=\textbf{Ret}^{max}.ret\_id\ \wedge\ \textbf{Ret}.ts=t^{\star}}\;\textbf{Ret}\Big).$
A.4 Treatment and control definitions
Products subject to SUP regulation are identified via keyword patterns $\mathcal{P}$ listed in Table 4:
$\textbf{Prod}^{SUP}$ $=\sigma_{\texttt{sup}(name)}(\textbf{Prod}^{+}),$ (12)
$\textbf{Prod}^{non}$ $=\pi_{\text{prod\_id}}(\textbf{Prod}^{+})\;-\;\pi_{\text{prod\_id}}(\textbf{Prod}^{SUP}),$
$\textbf{Prod}^{strict}$ $=\sigma_{\ name\ \texttt{ILIKE}\ \%\text{graphics card}\%}\big(\textbf{Prod}^{non}\ \bowtie\ \textbf{Prod}\big).$
Table 4: SUP keyword classes (illustrative ILIKE patterns).
Category Example patterns
Beverage cups %to-go becher%, %einwegbecher%, %kunststoffbecher%, %getränkebecher%
Food containers %takeaway box%, %kunststoffschale%
Films & wraps %verpackungsfolie%, %plastikfolie%
Plastic carrier bags %plastiktüte%, %einwegtragetasche%
Wet wipes %feuchttuch%, %hygienetuch%
Balloons %luftballon%, %partyballon%, %heliumballon%
A.5 Monthly aggregation and price index
Offers and clicks are aggregated to the product–retailer–month level:
Pbar $=\gamma_{\text{prod\_id},\ \text{ret\_id},\ \text{month}}\big(\mathrm{avg}(price)\rightarrow\overline{p}\big)(\textbf{Off}^{win}),$ (13)
Csum $=\gamma_{\text{prod\_id},\ \text{ret\_id},\ \text{month}}\big(\mathrm{sum}(clicks)\rightarrow clk\big)(\textbf{Clk}^{win}).$
Event time and three-month bins $b=\lfloor e/3\rfloor$ are assigned:
Pmon $=\mu_{\ e=\texttt{months\_between}(\text{month},m_{0}),\ b=\lfloor e/3\rfloor}(\textbf{Pbar}),$ (14)
Cmon $=\mu_{\ e=\texttt{months\_between}(\text{month},m_{0}),\ b=\lfloor e/3\rfloor}(\textbf{Csum}).$
Base prices at $m_{0}$ are used to construct a price index:
BaseP $=\pi_{\text{prod\_id},\ \text{ret\_id},\ \overline{p}\rightarrow\overline{p}_{0}}\ \sigma_{\text{month}=m_{0}}(\textbf{Pmon}),$ (15)
IndexAll $=\mu_{P=100\cdot\overline{p}/\overline{p}_{0}}\Big(\textbf{Pmon}\ \text{LOJ}_{(\text{prod\_id},\text{ret\_id})}\ \textbf{BaseP}\Big).$
A.6 Analysis samples
Clicks and retailer attributes are merged to form the analysis table:
$\textbf{ObsCore}\;:=\;\big(\textbf{IndexAll}\ \text{LOJ}_{(\text{prod\_id},\text{ret\_id},\text{month})}\ \textbf{Cmon}\big)\ \text{LOJ}_{\text{ret\_id}}\ \textbf{Ret}^{snap}.$ (16)
Treated and control samples are defined as:
$\textbf{Obs}^{T}$ $=\textbf{ObsCore}\ \bowtie_{\text{prod\_id}}\ \pi_{\text{prod\_id}}(\textbf{Prod}^{SUP}),$ (17)
$\textbf{Obs}^{C}$ $=\textbf{ObsCore}\ \bowtie_{\text{prod\_id}}\ \pi_{\text{prod\_id}}(\textbf{Prod}^{non}),$
$\textbf{Obs}^{S}$ $=\textbf{ObsCore}\ \bowtie_{\text{prod\_id}}\ \pi_{\text{prod\_id}}(\textbf{Prod}^{strict}).$
The variables exported to estimation are:
$\pi_{\ \text{prod\_id},\ \text{ret\_id},\ ret\_name,\ \text{month},\ e,\ b,\ P,\ \log P,\ clk\ }(\cdot).$ (18)
A.7 Estimation and interpretation
For each sample, the estimated regression is:
$Y_{ijt}\sim\sum_{b\neq 0}\mathbf{1}\{b_{t}=b\}\;\Big|\;\text{FE}_{i}+\text{FE}_{j},\qquad\text{clusters: }(i,j),$ (19)
implemented in R via fixest::feols(Y ~ i(event_bin, ref=0) | prod_id + ret_name, cluster ~ prod_id + ret_name). Figures report 90% confidence intervals for $Y=P$ (price index levels) and windowed DiD contrasts (6 months, 12 months, and full post–pre).
A.8 Timing and policy interpretation
Producer charges apply in Austria from 2022 onward, while Germany’s Single-Use Plastics Fund takes effect in 2024. Since $P=100\cdot\overline{p}/\overline{p}_{0}$ with base month February 2022, bin coefficients represent percentage-point deviations from that baseline. DiD contrasts summarize differences between average post- and pre-policy bins in these index points.